# references.bib

### v1

@@ 5.1 What Users Understand About How Their Data is Used
Another barrier to users’ ability to make informed decisions about their data is the use of deceptive design. These deceptive patterns (or “dark patterns”) are sets of design choices that coerce users to do things that aren’t in their best interest. Dark patterns can take the form of trick wording, comparison prevention, hidden costs, and nagging. These techniques can cause financial harm, loss of control (for instance over data usage), and discrimination [brignull-2023]. In the case of privacy policies as well, dark patterns may be employed to coerce users into making choices that are beneficial to the given service, but not necessarily beneficial to the user. Deceptive patterns are often used to make users part with more data than they may be aware of or be willing to disclose [nouwens2020]. Documenting and regulating dark patterns is challenging: they are deceptive, so users often are not aware of them, and users may also feel embarrassed about having been tricked, which complicates reporting [iobscura].
@@
It is possible to infer surprisingly invasive information from someone’s shopping habits; even to predict a consumer’s pregnancy before close family members are made aware [NYTprivsecrets] [Forbestargetpreg]. Time-stamped record of an individual’s cell phone location may be stored by mobile network operators for extended periods of time without the user’s knowledge, and can . [^34] Carpenter v. United States, 585 U.S. 296 (2018).
@@
Consumers tend to undervalue their privacy. This social phenomenon is captured by the “privacy paradox,” which identifies that people’s stated concerns about privacy are often not reflected in their online behaviors [privbehavecon]. Thus, giving consumers a choice to opt in or out will often not effectively translate their preferences. Additionally, the sheer volume and time constraints of decision-making with regards to privacy has been known to cause “consent fatigue” [murkyconsent]. Traditionally, consent has been provided in a binary format, like an on/off switch. However, as services offer wider ranges of settings and process increasingly more types of user data, privacy choices have become more granular. These consent requests compete for consumers limited time and attention, making it less likely for each request to be considered appropriately [WPprivpol].
@@ 5.2.1 Legal Consent
Legal consent presupposes a prior disclosure or legal notice. In digital services, legal notice may come in the form of a privacy statement, terms of service, or banner displayed by websites and apps. In contract law, privacy notices are regarded by a great majority of courts as a valid contracts, despite the extensive research (referenced in Section 5.1) showing that these notices are not read in practice [WurglerBarGill]. Terms of service and privacy policies are generally upheld as legally binding contracts even if users did not consent by an affirmative action like signing a document or checking a box labeled “I agree” while creating an account—American courts often treat a user’s continued use of a website or app as “implicit consent” to its terms and conditions.
@@ 5.2.3 Social Dimensions of Information Privacy
In his influential book Code and Other Laws of Cyberspace, Larry Lessig argued that laws are only one kind of the several regulations that affect the way people behave. [codeandcyberspace] Architecture (whether of buildings or code), markets, and social norms can regulate individuals’ behavior just as much, if not more than, law.
@@
Defining social norms precisely may be an impossible task in its own right. Helen Nissenbaum’s theory of contextual integrity argues that the term “privacy” must be understood depending on the context. [contextualintegrity] The norms around privacy in a health context may be completely different than in a school context, and from group to group. For example, if a child is diagnosed with ADHD by a private psychiatrist, it would be a severe violation of the child’s privacy for the psychiatrist to disclose that information outside of the child’s family. However, if the psychiatrist is employed by the child’s school, disclosure to the child’s teacher and the school’s accommodation staff may be expected. Blanket consent fails to account for these norms and undercuts privacy—addressing the flow of data itself is the only way to fully address privacy concerns.
@@
Linnet Taylor, Luciano Floridi, and Julie Cohen are three of many academics who argue for notions of “group privacy.” The motivating thought behind this concept is that giving each individual the last word on privacy concerning their personal data may not capture all the interests at stake. There are different flavors: Taylor and Floridi argue that informed consent can only be given by a data subject, so if the subject is a group, the consent of individuals is meaningless. [groupprivacy] The definition of the relevant group greatly affects the manner in which informed consent can be given. For example, acquiring the consent of a group chat would look very different than acquiring the consent of an ethnic group. Cohen argues that reduced privacy harms the normal functioning of society in a way that manifests far beyond the individual. [whatprivacyisfor] For example, individualized search engine results distort citizens’ perception of the world and make participation in the democratic process less likely to result in compromise. Both of these approaches acknowledge that an individual’s grant or withholding of consent is unlikely to capture the full set of issues involved.
@@
Several scholars have proposed changes to consent based on the idea of group privacy. August Bourgeus and Laurens Vandercruysse suggest a form of semi-automated consent, which would allow individuals to set their preferences accurately once and carry them throughout their online life. [reinvigoratingconsent] Daniel J. Solove advocates for “murky consent,” a baseline model which would rely on data holders accepting a duty of loyalty, duty to avoid thwarting reasonable expectations, and a duty to avoid unreasonable risk. [murkyconsent] Jack Balkin, and many scholars since, argues that holders of sensitive information should act as fiduciaries to data subjects [informationfiduciaries].f
@@ 6 Legal Implications of Integrating AI with E2EE
The integration of AI assistants into applications that process any form of E2EE data raises important considerations under current legal regimes. In this section, we begin with a survey of several E2EE messaging service providers’ Terms of Service and Privacy Policies to show that terms drafted by the service do not offer consumers meaningful choice in how their data is used. If a consumer aims to safeguard their data as much as possible, their only choice is to switch services. With this background in place, we overview relevant legal obligations in connection to data and consumer protection, considering how AI service providers are currently regulated in the US and in the EU. In tandem, we analyze how providers will likely be regulated if AI assistants and E2EE are integrated, highlighting aspects subject to ongoing regulation in the EU. Finally, we note regulatory trends across other jurisdictions and provide a brief note on how competition law may affect data-driven business operations.
@@ 6.1 Terms of Service and Privacy Policies of E2EE Services in Practice
As described in the prior section, when drafting terms of service and privacy policies, international companies must account for several significant regulations. In general, companies are incentivized to write overly broad terms of service that cover any possible future business development, because courts generally uphold broad contract terms, especially in the United States [Wurgler_Disclosure]. In the case of an E2EE messaging service, for example, a company might be incentivized to include a term in its fine print that says that it may stop offering E2EE at any time, even if all of its promotional material touts its E2EE, and even if it does currently offer strong E2EE and has no plans to cease providing it. Similarly, while E2EE messaging services may have high standards with regards to the third parties they contract with, they regularly attempt to disclaim liability for the actions of any other entity. Third party liability disclaimers are found in all four of the terms of service summarized below. [^38] This does not prevent companies from including quality assurances in contracts with vendors. In the event of a data breach, vendors can be liable to both direct customers and end consumers.[inreblackbaud] So, while E2EE messaging services may advertise lofty ideals, and even if they really do promote those ideals by deploying strong privacy technologies, their terms of service rarely make the same promises. WhatsApp’s terms of service would allow it to share user data for marketing purposes, a function that would conflict with the E2EE that it does, elsewhere, commit itself to providing. Such inconsistencies are common in terms of service because of the incentive to over-include.
Contractual privacy terms of select E2EE messaging services are summarized in Table 1. Note that the three services (WhatsApp, Facebook Messenger, and Telegram) that do not promise use of E2EE do not use it for messages with bots or businesses. Bots or businesses could store user messages in plaintext if they chose to—in other words, these three services have decided not to dictate third-party message handling practices, instead allowing the third-party’s privacy policy to govern these interactions. Finally, WhatsApp, Telegram, and Signal broadly market themselves as “secure” messaging platforms, and occasionally use the term “end-to-end encryption” in marketing. Facebook Messenger, on the other hand, does not mention security on its splash page, at the time of writing.
@@
Facebook Messenger No - spread across complex privacy center “On some Products, you can use end-to-end encrypted messages.” For marketing and to advertisers on Facebook, normal operations None
@@ 6.2 Overview of US Legal Considerations for E2EE & AI
As the sampled Terms of Service indicate, contract law is unlikely to create significant data protection to consumers in the absence of data protection and consumer protection laws.
@@
In the US, consumer protection law, a broad hammer mostly wielded by the FTC and state attorneys general, is likely to be provide stronger protections than data protection law, and has the flexibility to provide strong checks against the most egregious business practices.
@@ 6.2.1 Consumer Protection and the US Federal Trade Commission
Consumer protection law in the United States would apply to some implementations of AI in E2EE systems. At the federal level, the Federal Trade Commission Act of 1914 (15 U.S.C. § 45) created the Federal Trade Commission (FTC) and enabled it to investigate “unfair” or “deceptive” trade practices.
@@
Deceptive trade practices—the most relevant category here—are defined as those which constitute a “a misrepresentation, omission, or other practice, that misleads the consumer acting reasonably in the circumstances, to the consumer’s detriment.” [ftcdeceptionletter] The FTC investigates security-related incidents as deceptive practices. Such investigations are often prompted by data breaches and result in the FTC discovering misleading statements made to consumers about security. The FTC has discretion over which investigations to pursue—and must make choices, as it does not have the capacity to pursue every lead it encounters.
@@
For example, increased public use and scrutiny of Zoom in 2020 led to the discovery that Zoom was not protecting all video calls with the “end-to-end, 256-bit encryption” it advertised. The company settled with the FTC, agreeing constructively to implement a comprehensive security program and cease misrepresentations related to encryption [consentdecreezoom]. A dental software company advertising its use of “encryption,” unqualified, was found to be misleading customers because the level of encryption used did not meet industry standards [dentalsettlement]. A mortgage company settled with the FTC for advertising that the information it collected was encrypted, when in reality it was encrypted in transit but not once the company received it [mortgagesettlement].
In recent years, the FTC has decided to investigate dark patterns as deceptive practices as well [darkpatternslegalarticle]. In 2021 it published a policy statement against negative option marketing, and in 2022, it issued a report categorizing various dark patterns and the actions the agency would take against them.
@@
The FTC is not the only governmental agency capable of investigating deceptive practices. Many state attorneys general have similar powers granted by local versions of the FTC Act. [^39] For example, the New York State Attorney General investigates deceptive practices under the Deceptive Practices Act, N.Y. Gen. Bus. Law § 349. State attorneys general are important consumer protection enforcers because they can address smaller or more local concerns that the FTC would miss or lack the bandwidth to address. They can also be significant regulators in their own right, and collaborate between states to take on the most sophisticated corporations.
Under FTC precedent, there are clear lines that an E2EE messaging app could not cross without being found deceptive. If the app is marketed as using “end-to-end encryption,” established FTC precedents demonstrate that the app must use end-to-end encryption throughout the processing of the relevant data [consentdecreezoom, mortgagesettlement], have E2EE on by default [consentdecreezoom], and meet industry standards [dentalsettlement].
@@ 6.3.1 EU Legal Background
In many ways, legal obligations for data-driven services in the EU extend beyond the disclosure and consent model. For example, the GDPR sets out key principles that impose additional constraints upon collecting, storing, and using personal data. [^40] For example, the “lawfulness, fairness and transparency”, “purpose limitation” and “data minimization” principles are important guardrails enshrined in the GDPR. Regulation (EU) 2016/679 (“GDPR”), Article 5(1). Under the GDPR, obtaining consent is not the only basis to legitimize the processing of user data—it is just one among six. Many user data-related activities within companies can rely on the broad alternative basis of “legitimate interests,” that is, where processing is necessary for a legitimate interest pursued by the organization. [^41] Regulation (EU) 2016/679 (“GDPR”), Article 6(1)(f). For instance, analyzing customer data for fraud prevention and signs of criminal activity or engaging in direct marketing with preexisting clients are typically deemed activities within a company’s legitimate interests [icolegint]. The legal basis of “legitimate interests” is often seen by organizations as more flexible and less cumbersome than obtaining user consent. [^42] Despite the requirement to ensure that organizational interests do not override users’ fundamental rights (known as the “balancing test”). Regulation (EU) 2016/679 (“GDPR”), Article 6(1)(f).
@@
This is because the GDPR sets higher standards for user consent: it must be freely given, specific, informed, unambiguous, and may be withdrawn (or opted out) at any time. Employing dark patterns to obtain user consent or to hide mechanisms to opt out may violate the GDPR and other overlapping regulations, such as the 2022 Digital Services Act (DSA) and the Unfair Commercial Practices Directive (“UCP” Directive)[ECconsumanip]. The European Data Protection Board (EDPB) provided examples of deceptive techniques used by social media platforms that do not qualify as “freely-given” consent [edpbdarkpat]. These include “forced” consent, where refusing to consent leads to a denial of the service, and “bundled” consent, when consent is bundled with the acceptance of the platform’s terms and conditions.
@@
The GDPR is not the only data protection law in the EU. Other data protection laws do mandate obtaining user consent in certain situations. For example, the surveillance of private communications necessitates user consent. [^43] Directive (EU) 2002/58 on Privacy and Electronic Communications (“ePrivacy Directive”), Article 5(1). Another example is the storing of non-essential browser cookies and other tracking technologies in user devices—a requirement that has reshaped the appearance of websites with the display of consent-collecting “cookie banners”. [^44] Directive (EU) 2002/58 on Privacy and Electronic Communications (“ePrivacy Directive”), Article 5(3). The requirement to obtain user consent in these instances is mandated by the ePrivacy Directive (Directive (EU) 2002/58 on Privacy and Electronic Communications). Article 5 of the ePrivacy Directive establishes a general prohibition on storing information or gaining access to stored information on users’ end devices without obtaining prior consent, including listening, tapping, or other kinds of interception or surveillance of communications and related metadata. This legislation was enacted to complement the 1995 Directive on Data Protection, the GDPR’s predecessor, to secure the confidentiality of electronic communications in protection of “the right to respect for private and family life”, as set forth in the EU’s Charter of Fundamental Rights. [^45] Charter of Fundamental Rights of the European Union (2000), Article 7.
@@
While the ePrivacy Directive clearly applies to providers of network services, it is less clear whether it applies to “over-the-top” (OTT) communication services such as WhatsApp, Telegram and Zoom. In 2017, the European Commission adopted a proposal to replace the Directive with a new ePrivacy Regulation in an effort to update the law and explicitly cover OTT services [EPeprivacy]. Attempts to advance the reform received significant pushback from industry players and were undercut by the passing of a temporary derogation that allows the scanning of private communications to detect online Child Sexual Abuse Material [edriepriv]. As of May 2021, the draft ePrivacy Regulation still faces complicated trilogue negotiations in the EU Parliament and is currently blocked [EPeprivacy]. Despite the uncertainty facing the reform proposal, the European Data Protection Board (EDPB) recommends that a new ePrivacy Regulation must under no circumstances lower the level of protection offered by the current ePrivacy Directive, as the confidentiality of electronic communications is a fundamental right that requires specific protection [edpbepriv]. The EDPB further recommends that any exceptions must be narrowly tailored, and should not lead to the systematic monitoring of electronic communication content, nor allow providers or third parties to circumvent any encryption.
@@
The EU’s recently enacted AI Act reinforces the applicability of both the ePrivacy Directive and the GDPR to AI systems, highlighting that these laws establish the basis for sustainable and responsible data processing. [^46] Regulation (EU) 2024/1689 (“EU AI Act”), Recital 10 and Article 2(7). Recital 69 of the AI Act notes that the right to privacy must be guaranteed throughout the entire lifecycle of the AI system, and that measures taken by providers to ensure compliance may include not only anonymization and encryption, but also “the use of technology that permits algorithms to be brought to the data and allows training of AI systems without the transmission between parties or copying of the raw or structured data themselves”.
6.3.2 Training AI in the EU: Ongoing Controversies

### v2

@@ 5 Background II: Consent, Privacy, Users, and Areas of Law
Next, we turn to legal and normative considerations around clear disclosures about and meaningful consent to the use of AI integrated with E2EE systems, complementing our preceding analysis of when and how such systems can be technically compatible with the confidentiality promised by E2EE. The legal landscape around E2EE and AI is rapidly evolving, raising questions about user protection and data practices that, in turn, have implications on security and confidentiality. This section provides preliminary background on data practices (Section 5.1), notice and consent (Section 5.2), and relevant areas of law (Section 5.3).
@@ 5.1 What Users Understand About How Their Data is Used
Another barrier to users’ ability to make informed decisions about their data is the use of deceptive design. These deceptive patterns (or “dark patterns”) are sets of design choices that coerce users to do things that aren’t in their best interest. Dark patterns can take the form of trick wording, comparison prevention, hidden costs, and nagging. These techniques can cause financial harm, loss of control (for instance over data usage), and discrimination [brignull-2023]. In the case of privacy policies, dark patterns may be employed to coerce users into making choices that are beneficial to the given service, but not necessarily beneficial to the user. Deceptive patterns are often used to make users part with more data than they may be aware of or be willing to disclose [nouwens2020]. Documenting and regulating dark patterns is challenging: they are deceptive, so users often are not aware of them, and users may also feel embarrassed about having been tricked, which complicates reporting [iobscura].
@@
It is possible to infer surprisingly invasive information from someone’s shopping habits; even to predict a consumer’s pregnancy before close family members are made aware [NYTprivsecrets] [Forbestargetpreg]. Time-stamped record of an individual’s cell phone location may be stored by mobile network operators for extended periods of time without the user’s knowledge. [^43] Carpenter v. United States, 585 U.S. 296 (2018).
@@
Consumer actions tend to undervalue their privacy. This social phenomenon is captured by the “privacy paradox,” which identifies that people’s stated concerns about privacy are often not reflected in their online behaviors [privbehavecon]. Thus, giving consumers a choice to opt in or out will often not effectively translate their preferences. Additionally, the sheer volume and time constraints of decision-making with regards to privacy has been known to cause “consent fatigue” [murkyconsent]. Traditionally, consent has been provided in a binary format, like an on/off switch. However, as services offer wider ranges of settings and process increasingly more types of user data, privacy choices have become more granular. These consent requests compete for consumers limited time and attention, making it less likely for each request to be considered appropriately [WPprivpol].
@@ 5.2.1 Legal Consent
Legal consent presupposes a prior disclosure or legal notice. In digital services, legal notice may come in the form of a privacy statement, terms of service, or banner displayed by websites and apps. In contract law, privacy notices are regarded by a great majority of courts as a valid contracts, despite extensive research (see Section 5.1) showing that these notices are not read in practice [WurglerBarGill]. Terms of service and privacy policies are generally upheld as legally binding contracts even if users did not consent by an affirmative action like signing a document or checking a box labeled “I agree” while creating an account—American courts often treat a user’s continued use of a website or app as “implicit consent” to its terms and conditions.
@@ 5.2.3 Social Dimensions of Information Privacy
In his influential book Code and Other Laws of Cyberspace, Larry Lessig argued that laws are only one kind of the several regulations that affect the way people behave [codeandcyberspace]. Architecture (whether of buildings or code), markets, and social norms can regulate individuals’ behavior just as much, if not more than, law.
@@
Defining social norms precisely may be an impossible task in its own right. Helen Nissenbaum’s theory of contextual integrity argues that the term “privacy” must be understood depending on the context [contextualintegrity]. The norms around privacy in a health context may be completely different than in a school context, and from group to group. For example, if a child is diagnosed with ADHD by a private psychiatrist, it would be a severe violation of the child’s privacy for the psychiatrist to disclose that information outside of the child’s family. However, if the psychiatrist is employed by the child’s school, disclosure to the child’s teacher and the school’s accommodation staff may be expected. Blanket consent fails to account for these norms and undercuts privacy—addressing the flow of data itself is the only way to fully address privacy concerns.
@@
Linnet Taylor, Luciano Floridi, and Julie Cohen are three of many academics who argue for notions of “group privacy.” The motivating thought behind this concept is that giving each individual the last word on privacy concerning their personal data may not capture all the interests at stake. There are different flavors: Taylor and Floridi argue that informed consent can only be given by a data subject, so if the subject is a group, the consent of individuals is meaningless [groupprivacy]. The definition of the relevant group greatly affects the manner in which informed consent can be given. For example, acquiring the consent of a group chat would look very different than acquiring the consent of an ethnic group. Cohen argues that reduced privacy harms the normal functioning of society in a way that manifests far beyond the individual [whatprivacyisfor]. For example, individualized search engine results distort citizens’ perception of the world and make participation in the democratic process less likely to result in compromise. Both of these approaches acknowledge that an individual’s grant or withholding of consent is unlikely to capture the full set of issues involved.
@@
Several scholars have proposed changes to consent based on the idea of group privacy. August Bourgeus and Laurens Vandercruysse suggest a form of semi-automated consent, which would allow individuals to set their preferences accurately once and carry them throughout their online life [reinvigoratingconsent]. Daniel J. Solove advocates for “murky consent,” a baseline model which would rely on data holders accepting a duty of loyalty, duty to avoid thwarting reasonable expectations, and a duty to avoid unreasonable risk [murkyconsent]. Jack Balkin, and many scholars since, argues that holders of sensitive information should act as fiduciaries to data subjects [informationfiduciaries].
@@ 6 Legal Implications of Integrating AI with E2EE
The integration of AI assistants into applications that process any form of E2EE data raises important considerations under current legal regimes, which are related to the security considerations we have laid out thus far, as well as broader legal issues around disclosure and consent.
In this section, we begin with a brief survey of privacy terms in several E2EE messaging service providers’ Terms of Service and Privacy Policies (Section 6.1). We then overview relevant legal obligations under data protection and consumer protection law, considering how AI service providers are currently regulated in the US (Section 6.2) and in the EU (Section 6.3). The specific issue of AI processing of E2EE data has not yet come up in legal proceedings, but our discussion provides background on relevant existing law, as well as ongoing controversies and uncertainties, that will inform how AI processing of E2EE data is treated under US and EU law. Finally, we note regulatory trends across other jurisdictions (Section 6.4) and provide a brief note on how competition law may affect data-driven business operations (Section 6.5).
@@ 6.1 Terms of Service and Privacy Policies of E2EE Services in Practice
When drafting terms of service and privacy policies, international companies must account for several significant regulations. In general, companies are incentivized to write overly broad terms of service that cover any possible future business development, because courts generally uphold broad contract terms, especially in the United States [Wurgler_Disclosure]. In the case of an E2EE messaging service, for example, a company might be incentivized to include a term in its fine print that says that it may stop offering E2EE at any time, even if all of its promotional material touts its E2EE, and even if it does currently offer strong E2EE and has no plans to cease providing it. Similarly, while E2EE messaging services may have high standards with regards to the third parties they contract with, they regularly attempt to disclaim liability for the actions of any other entity. Third party liability disclaimers are found in all four of the terms of service summarized below. [^47] This does not prevent companies from including quality assurances in contracts with vendors. In the event of a data breach, vendors can be liable to both direct customers and end consumers [inreblackbaud]. So, while E2EE messaging services may advertise lofty ideals, and even if they really do promote those ideals by deploying strong privacy technologies, sometimes their terms of service may not make the same promises. For example, WhatsApp’s terms of service would allow it to share user data for marketing purposes, a function that would conflict with the E2EE that it does, elsewhere, commit itself to providing. Such inconsistencies are common in terms of service because of the incentive to over-include.
The contractual privacy terms of four messaging services which offer E2EE features are summarized in Table 2. Notably, the three services (WhatsApp, Meta’s Messenger, and Telegram) that do not promise use of E2EE do not use it for messages with bots or businesses. Bots or businesses could store user messages in plaintext if they chose to—in other words, these three services have decided not to dictate third-party message handling practices, instead allowing the third-party’s privacy policy to govern these interactions. Finally, WhatsApp, Telegram, and Signal broadly market themselves as “secure” messaging platforms, and occasionally use the term “end-to-end encryption” in marketing. Meta’s Messenger, on the other hand, does not mention security on its splash page, at the time of writing.
Of the four services discussed, WhatsApp’s privacy policy notably promises less security than its marketing materials promise, and than its product actually provides. To our knowledge, the service as provided at the time of writing falls into Section 3.2’s Category 2, consistent with marketing materials, while the privacy policy describes a Category 3 system.
@@
Meta’s Messenger No - spread across complex privacy center “On some Products, you can use end-to-end encrypted messages.” For marketing and to advertisers on Facebook, normal operations None
@@ 6.2 Overview of US Legal Considerations for E2EE & AI
In the US, consumer protection law, a broad hammer mostly wielded by the FTC and state attorneys general, is likely to be provide stronger protections than data protection law, and has the flexibility to provide strong checks against the most egregious business practices, as discussed next.
@@ 6.2.1 Consumer Protection and the US Federal Trade Commission
Consumer protection law in the United States would constrain some implementations of AI in E2EE systems. At the federal level, the Federal Trade Commission Act of 1914 (15 U.S.C. § 45) created the Federal Trade Commission (FTC) and enabled it to investigate “unfair” or “deceptive” trade practices.
@@
Deceptive trade practices—the most relevant category here—are defined as those which constitute “a misrepresentation, omission, or other practice, that misleads the consumer acting reasonably in the circumstances, to the consumer’s detriment” [ftcdeceptionletter]. The FTC investigates security-related incidents as deceptive practices. Such investigations are often prompted by data breaches and result in the FTC discovering misleading statements made to consumers about security. The FTC has discretion over which investigations to pursue—and must make choices, as it does not have the capacity to pursue every lead it encounters.
@@
For example, increased public use and scrutiny of Zoom in 2020 led to the discovery that Zoom was not protecting all video calls with the “end-to-end, 256-bit encryption” it advertised. The company settled with the FTC, agreeing constructively to implement a comprehensive security program and cease misrepresentations related to encryption [consentdecreezoom]. A dental software company advertising its use of “encryption,” unqualified, was found in an FTC proceeding to be misleading customers because the level of encryption used did not meet industry standards [dentalsettlement]. A mortgage company settled with the FTC for advertising that the data it collected was encrypted, when in reality it was encrypted in transit but not after receipt [mortgagesettlement].
There is no precise technical specification that companies must meet to avoid being prosecuted for deceptive practices. [^48] But there is a roughly defined floor, below which companies can be investigated for unfair practices. This comes up often when companies are subjects of a data breach compromising unencrypted credit card, passport, or social security numbers [mariottstarwoodbreach]. That said, under FTC precedent, there are clear lines that an E2EE messaging app could not cross without being found deceptive. If an app is marketed as using “end-to-end encryption,” established FTC precedents demonstrate that the app must use end-to-end encryption throughout the processing of the relevant data [consentdecreezoom, mortgagesettlement], have E2EE on by default [consentdecreezoom], and meet industry standards [dentalsettlement]. In other words, the marketing of a service in Category 5 (No E2EE) defined in Section 3.2 with an unqualified claim that the product uses “end-to-end encryption” would be considered deceptive under FTC precedent. We stress, again, that the FTC has discretion over which cases it pursues and prioritizes—so despite the above, an app in violation of the above would not necessarily be investigated or suffer consequences.
@@
The FTC is not the only governmental agency capable of investigating deceptive practices. Many state attorneys general have similar powers granted by local versions of the FTC Act. [^49] For example, the New York State Attorney General investigates deceptive practices under the Deceptive Practices Act, N.Y. Gen. Bus. Law § 349. State attorneys general are important consumer protection enforcers because they can address smaller or more local concerns that the FTC would miss or lack the bandwidth to address. They can also be significant regulators in their own right, and collaborate between states to take on the most sophisticated corporations. With uncertainty around the enforcement objectives of the FTC during the second Trump administration [ferguson-ftc], state attorneys general are likely to take on even more importance as enforcers of consumer protection laws.
@@ 6.3.1 EU Legal Background
In many ways, legal obligations for data-driven services in the EU extend beyond the disclosure and consent model described above. Next, we overview key relevant EU legislation.
GDPR
@@
The GDPR sets out key principles that impose additional constraints upon collecting, storing, and using personal data. [^50] For example, the “lawfulness, fairness and transparency”, “purpose limitation” and “data minimization” principles are important guardrails enshrined in the GDPR. Regulation (EU) 2016/679 (“GDPR”), Article 5(1). Under the GDPR, obtaining consent is not the only basis to legitimize the processing of user data—it is just one among six. Many user data-related activities within companies can rely on the broad alternative basis of “legitimate interests,” that is, where processing is necessary for a “legitimate interest” pursued by the organization. [^51] Regulation (EU) 2016/679 (“GDPR”), Article 6(1)(f). For instance, analyzing customer data for fraud prevention and signs of criminal activity or engaging in direct marketing with preexisting clients are typically deemed activities within a company’s legitimate interests [icolegint]. The legal basis of “legitimate interests” is often seen by organizations as more flexible and less cumbersome than obtaining user consent. [^52] Despite the requirement to ensure that organizational interests do not override users’ fundamental rights (known as the “balancing test”). Regulation (EU) 2016/679 (“GDPR”), Article 6(1)(f).
@@
This is because the GDPR sets higher standards for user consent: it must be “freely given,” “specific,” “informed,” “unambiguous,” and may be withdrawn (or opted out) at any time. Employing dark patterns to obtain user consent or to hide mechanisms to opt out may violate the GDPR and other overlapping regulations, such as the 2022 Digital Services Act (DSA) and the Unfair Commercial Practices Directive (“UCP” Directive) [ECconsumanip]. The European Data Protection Board (EDPB) provided examples of deceptive techniques used by social media platforms that do not qualify as “freely-given” consent [edpbdarkpat]. These include “forced” consent, where refusing to consent leads to a denial of the service, and “bundled” consent, when consent is bundled with the acceptance of the platform’s terms and conditions. These heightened protections are valuable; at the same time, the GDPR still offers an imperfect regime for protecting user rights and falls short of the ideals of freely given, unambiguous, informed consent based on mutual assent (e.g., [cofoneconsent]).
The GDPR is just one piece of data protection law in the EU. Other data protection legislation does mandate obtaining user consent in certain situations, as discussed next.
ePrivacy Directive
@@
The ePrivacy Directive (Directive (EU) 2002/58 on Privacy and Electronic Communications) requires obtaining user consent for the surveillance of private communications [^53] Directive (EU) 2002/58 on Privacy and Electronic Communications (“ePrivacy Directive”), Article 5(1). and the storing of non-essential browser cookies and other tracking technologies in user devices—a requirement that has reshaped the appearance of websites with the display of consent-collecting “cookie banners”. [^54] Directive (EU) 2002/58 on Privacy and Electronic Communications (“ePrivacy Directive”), Article 5(3). Article 5 of the ePrivacy Directive establishes a general prohibition on storing information or gaining access to stored information on users’ end devices without obtaining prior consent, including listening, tapping, or other kinds of interception or surveillance of communications and related metadata.
@@
As the ePrivacy Directive was originally designed to regulate traditional telecom operators, it is not completely clear whether it applies to “over-the-top” (OTT) communication services such as WhatsApp, Telegram and Zoom. In 2017, the European Commission adopted a proposal to replace the Directive with an updated ePrivacy Regulation that would explicitly cover OTT services [EPeprivacy]. Attempts to advance the reform received significant pushback from industry players, and were undercut by the passing of a temporary derogation that allows the scanning of private communications to detect online Child Sexual Abuse Material [edriepriv]. After being blocked for several years due to complicated trilogue negotiations in the EU Parliament, the ePrivacy Regulation proposal was withdrawn by the EU Commission in February 2025 [EPeprivacy] [EU-privreg-ditched].
Despite the uncertainty facing potential reforms, the European Data Protection Board (EDPB) recommends that a new ePrivacy Regulation must under no circumstances lower the level of protection offered by the current ePrivacy Directive, as the confidentiality of electronic communications is a fundamental right that requires specific protection [edpbepriv]. The EDPB further recommends that any exceptions must be narrowly tailored, and should not lead to the systematic monitoring of electronic communication content, nor allow providers or third parties to circumvent any encryption.
AI Act
@@
The EU’s recently enacted AI Act reinforces the applicability of both the ePrivacy Directive and the GDPR to AI systems, highlighting that these laws establish the basis for sustainable and responsible data processing. [^55] Regulation (EU) 2024/1689 (“EU AI Act”), Recital 10 and Article 2(7). Recital 69 of the AI Act notes that the rights to privacy and to protection of personal data—which are fundamental rights in the EU [^56] Charter of Fundamental Rights of the European Union (2000), Articles 7 and 8. —must be “guaranteed throughout the entire lifecycle of [an] AI system,” and that measures taken by providers to ensure compliance may include anonymization, encryption, and “the use of technology that permits algorithms to be brought to the data and allows training of AI systems without the transmission between parties or copying of the raw or structured data themselves”.
6.3.2 Training AI in the EU and the UK: Ongoing Controversies